# Expedited Noise Spectroscopy of Transmon Qubits

### v1

@@ IV.1 Network Performance on the Test Data
We trained the neural network on a subset of the generated pairs of decoherence curves $C(t)$ and noise spectra $S(\omega)$ (the ‘training’ set). To then test the ability of the trained network to produce the correct noise spectrum from a single decoherence curve, we used another subset of the data (the ‘test’ set) and compared the noise spectrum produced by our network with the originally generated one. The solid lines in Fig. 1(c) represent the ‘test’ noise spectra, which were used to obtain the corresponding ‘test’ coherence curves, shown in solid lines in Fig. 1(a), by applying Eq. 3. The dashed lines in Fig. 1(c) are then the noise spectra obtained by applying the trained model to the ‘test’ coherence curves. The accuracy of the model prediction is excellent with a mean absolute error of 3.6%, as shown in the inset. Note that this error is largely limited by the random noise added to the training data to mimic the experimental signal-to-noise ratio, suggesting that the performance of the network is near optimal. As a final check, the neural net-produced noise spectra were used to obtain decoherence curves to see how well those match the ‘test’ decoherence curves that had been given as an input to the network. These are plotted in dashed lines in Fig. 1(a), and indeed show good agreement with the ‘test’ decoherence curves, while demonstrating the de-noising capabilities of our network.
IV.2 Network Performance on IBM Qubits
@@
IV.3 Optimization of Dynamical Decoupling Pulse Sequence
@@
The predicted noise spectra were then used to develop the optimal sequences with varied number $n$ of $\pi$-pulses. Based on the $1/\omega$ frequency cut-off in the noise spectra, we estimated that the $n=8$ $\pi$-pulse sequences should provide the maximum advantage. Our experimental results, obtained on IBM Osaka quantum processor for two different qubits are displayed in Fig. 3(a) and (e) with the underlying noise spectra plotted in the insets. The results reveal two notable outcomes. Firstly, the customized pulse sequences shown in Fig. 3(b) and (f) do not show noticeable improvement as compared to traditional protocols displayed in Fig.3(c) and (g). This is due to the fact that the noise spectra are heavily white noise dominated and, consequently, the modified filter functions shown in Fig.3(d) and (h) are unable to suppress the overlap. Secondly, despite implementing non-trivial bespoke pulse sequences, the experimental decoherence curves closely match with those predicted via the optimization algorithm. This authenticates the accuracy of the noise spectra predicted by the network. The overall methodology is, thus, a step forward in the noise-adapted optimal quantum error mitigation techniques. The effectiveness of the optimized sequences should certainly be notable for other qubit systems where white noise is not the leading type of the environmental noise.
IV.4 Rapid Time Dependent Noise Spectroscopy
@@
In order to capture the time-evolution of the noise around IBM qubits, we performed time-dependent noise spectroscopy experiments on a qubit. To achieve this, ten $T_{1}$, and ten $T_{2}$ measurements were performed consecutively while keeping the experimental parameters unchanged. Note that each run consists of 4000 shots to attain sufficient signal-to-noise ratio which required approximately 10 minutes to complete. The results displayed in Fig.4(a) and (b) demonstrate that we are indeed able to record the rapid changes in $T_{1}$ and $T_{2}$ curves. Moreover, the extracted noise spectra presented in Fig.4(c) indicate that the discontinuous jumps in the dephasing and relaxation data are neither time-correlated nor entirely interdependent. These observations confirm that our noise spectroscopy model is a very effective tool to extract quasi-instantaneous noise profiles.
@@ V OUTLOOK
Owing to the universality of the technique discussed here, we anticipate this study will help examine and benchmark various quantum systems in the near future. We expect that the noise characterization methodology discussed here will provide useful insights to improve the engineering aspects of quantum systems. The fidelity of the current methodology can be improved by taking into account advanced gates Baum et al. (2021), or DRAG pulses Werninghaus et al. (2021); Hyyppä et al. (2024), which are optimized to reduce leakage error, instead of square pulses. This should be possible owing to the recent advancements in the filter function formalism, which allows for analytical representation of arbitrary quantum control sequence Soare et al. (2014); Cerfontaine et al. (2021); Chalermpusitarak et al. (2021); Le et al. (2022). Moreover, if this methodology is employed in conjunction with appropriate instrumentation such as Field Programmable Gate Arrays (FPGAs) Goldblatt et al. (2024), then instantaneous error mitigation strategies can potentially be developed.
It is important to reemphasize that the decoherence functional, $\chi(t)$, which depends on the knowledge of the environmental noise spectrum is intricately linked to the error rates of the various noise operators in the Kraus picture. Since Kraus formalism is commonly employed in the design and implementation of quantum error correction (QEC) protocols, having precise knowledge of the decoherence functional opens up new avenues for noise-adapted error correction protocols Biswas et al. (2024). Indeed, QEC protocols can then be tailored to deal with a specific spectral range of noise. With the help of a trained model, time-dependent quantum error correction protocols can become a foreseeable reality.
@@ VI Acknowledgements
S.D. thanks the Indian Institute of Technology, Madras, India, for its seed and initiation grants. We acknowledge the use of facilities supported by a grant from the Mphasis F1 Foundation given to the Centre for Quantum Information, Communication, and Computing (CQuICC). Finally, we appreciate the use of IBM Quantum for this work. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team.
@@ Appendix A Noise Model: Kraus Operator Picture
where the physical process map $\mathcal{E}$ on the system operator Hilbert space $\mathcal{H}_{S}$ must be completely positive (CP) The complete positivity of the map implies the existence of an operator—sum decomposition for the map Nielsen and Chuang (2011), of the following form
@@
We will now describe two examples of such CPTP maps – also referred to as quantum channels Nielsen and Chuang (2011)— which become important in the context of superconducting qubits. The first is the so called amplitude damping channel, which characterizes the effects due to loss of energy from a quantum system. Specifically, it describes energy dissipation in a two-level system. Let $|0\rangle$ denote the ground state and $|1\rangle$ some excited state of a qubit. Then, the amplitude damping channel denoted as $\mathcal{E}^{AD}=\{E_{0}^{AD},E_{1}^{AD}\}$, is described by the following pair of Kraus operators Nielsen and Chuang (2011).
@@
$=\begin{pmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{pmatrix}$
@@
$E_{0}^{PD}$ $=\sqrt{p}I=\begin{pmatrix}\sqrt{p}&0\\ 0&\sqrt{p}\end{pmatrix}$
@@
$\rho=\begin{pmatrix}|\alpha|^{2}&\alpha\beta^{*}\\ \alpha^{*}\beta&|\beta|^{2}\end{pmatrix}\xrightarrow{noise}\rho_{noisy}=\mathcal{E}^{AD}\circ\mathcal{E}^{PD}(\rho)$
$=\begin{pmatrix}1+(|\alpha|^{2}-1)e^{-\Gamma_{1}t}&\alpha\beta^{*}e^{-\frac{\Gamma_{1}}{2}t}e^{-\Gamma_{\phi}t}\\ \alpha^{*}\beta e^{-\frac{\Gamma_{1}}{2}t}e^{-\Gamma_{\phi}t}&|\beta|^{2}e^{-\Gamma_{1}t}\end{pmatrix}$ (8)
@@ Appendix B Neural Network
We used TensorFlow Keras Python module to train our convolutional neural network and to perform the noise spectra predictions from the experimentally obtained decoherence curves. The autoencoder-type network, as described in Table 2 alternates between convolutional and pooling operations to progressively extract features from the input data, and the upsampling restores the data’s original dimensions while enhancing feature representation. The final dense layer aims to produce a linear output appropriate for regression or other continuous output tasks. Our network is trained for approximately 75 epochs, achieving an accuracy loss lower than 4%. The training and validation losses, measured as mean absolute error, are monitored over epochs. Furthermore, to obtain a high degree of prediction accuracy, it is essential to constrain the training data appropriately. We fit the generated decoherence curves using a stretched exponential function:
$\tilde{C}(t)=e^{-(t/T_{\phi})^{p}}$ (9)
Although, the simulated curves cannot be exactly fitted with Eq.9, this method provides a good measure to filter the training data. Fig 5 shows the final distribution of the stretching factor $p$, and of inverse decay rate $T_{\phi}$ in the training dataset. All the experimental curves lie well within the chosen bounds.
Figure 5: Distribution of fitted $p$ and $T_{\phi}$ in the training dataset

### v2

@@ IV.1 Network Performance on the Test Data
Figure 1: (a) Randomly chosen 5 test input decoherence curves as functions of evolution time; the dashed lines denote the curves evaluated from the predicted $S(\omega)$. (b) A schematic diagram of the neural network developed in this work. (c) The numerically generated (solid lines) and the predicted (dashed line) noise spectra associated with the input decoherence curves. The inset shows a histogram of the estimation errors in coherence curves.
We trained the neural network on a subset of the generated pairs of decoherence curves $C(t)$ and noise spectra $S(\omega)$ (the ‘training’ set). To then test the ability of the trained network to produce the correct noise spectrum from a single decoherence curve, we used another subset of the data (the ‘test’ set) and compared the noise spectrum produced by our network with the originally generated one. The solid lines in Fig. 1(c) represent the ‘test’ noise spectra, which were used to obtain the corresponding ‘test’ coherence curves, shown in solid lines in Fig. 1(a), by applying Eq. 3. The dashed lines in Fig. 1(c) are then the noise spectra obtained by applying the trained model to the ‘test’ coherence curves. As a final check, the neural net-produced noise spectra were used to obtain decoherence curves to see how well those match the ‘test’ decoherence curves that had been given as an input to the network. These are plotted in dashed lines in Fig. 1(a), and indeed show good agreement with the ‘test’ decoherence curves, while demonstrating the de-noising capabilities of our network. The accuracy of the model prediction is excellent with a mean absolute error of 3.6%, as shown in the inset. Note that this error is largely limited by the random noise added to the training data to mimic the experimental signal-to-noise ratio, suggesting that the performance of the network is near optimal
IV.2 Error Analysis and Robustness of Neural Network-Based Noise Spectroscopy
Figure 2: The heatmap displaying percentage mode values associated with the prediction errors. The $x$-axis denotes the percentage Gaussian noise used while training the model, while the $y$-axis represents the percentage Gaussian noise employed during the testing phase. The histograms at the bottom show the distribution of the percentage errors in the individual $S(\omega)$ predictions for the two scenarios specified in the heatmap.
The neural network-based approach to noise spectroscopy offers significant benefits over traditional methods, particularly for cases where the signal-to-noise ratio is relatively low due to a high-speed measurement protocol with a small number of shots. To demonstrate this advantage, we performed a systematic analysis of how the prediction errors scale with the unavoidable experimental noise. We trained five different networks using the same underlying training data, but with varying levels of synthetic experimental noise (i.e. Gaussian noise) ranging from 1% to 9%. We then used the test data having different degrees of synthetic experimental noise added to them to obtain predicted noise spectra for each of these combinations. It is important to note that in the shot-noise dominated regime, these experimental noise levels correspond to an 81-fold change in the experimental duration. Subsequently, we analyze the distribution of the errors for each case to extract the mode value; the individual mode values are depicted in the heatmap shown in Fig. 2, whereas the histograms of the errors for the two specific cases highlighted in the heatmap are shown in the bottom panel. The analysis supports the following major advantages of the neural network based approach -
@@ IV.2 Error Analysis and Robustness of Neural Network-Based Noise Spectroscopy
  1. 1.
    First, the neural network trained with intermediate levels of experimental noise, e.g. $3\%$ noise, is extremely resilient to the changes in signal-to-noise ratio. This signifies that the experimenter can reduce the measurement time significantly without a substantial reduction in prediction accuracy. Moreover, this also means that the experimenter need not precisely replicate the levels of experimental noise used during training. This robustness against the noise is a characteristic feature of convolutional neural networks, which are often used for image denoising.
  2. 2.
    Secondly, classical optimizers are extremely sensitive to initial guesses and frequently get trapped at local minima. On the contrary, it is known that over-parametrized neural networks tend to obtain global minima with high likelihood Choromanska et al. (2015). Histograms with a well-defined maximum observed in the case of the neural network (Fig. 2) as opposed to a distribution that appears to be reasonably random in the case of classical optimizers (see Fig. 3 in the Supplementary Material) indeed validate this claim.
  3. 3.
    Finally, the prediction speed of neural networks is undeniably superior to that of classical optimizers. Classical optimizers are known to exhibit sluggish performance and convergence issues, especially in the context of multivariate functions.
Building on these validations, we carried out neural network–enabled quantum noise spectroscopy on a physical quantum device as discussed in the following section.
IV.3 Network Performance on IBM Qubits
@@
IV.4 Optimization of Dynamical Decoupling Pulse Sequence
@@
The predicted noise spectra were then used to develop the optimal sequences with varied number $n$ of $\pi$-pulses. Based on the $1/\omega$ frequency cut-off in the noise spectra, we estimated that the $n=8$ $\pi$-pulse sequences should provide the maximum advantage (see SI Section 3 and the figure therein). Our experimental results, obtained on IBM Osaka quantum processor for two different qubits are displayed in Fig. 4(a) and (e) with the underlying noise spectra plotted in the insets. The results reveal two notable outcomes. Firstly, the customized pulse sequences shown in Fig. 4(b) and (f) do not show noticeable improvement as compared to traditional protocols displayed in Fig.4(c) and (g). This is due to the fact that the noise spectra are heavily white noise dominated and, consequently, the modified filter functions shown in Fig.4(d) and (h) are unable to suppress the overlap. Secondly, despite implementing non-trivial pulse sequences, the experimental decoherence curves closely match with those predicted via the optimization algorithm. This authenticates the accuracy of the noise spectra predicted by the network. The overall methodology is, thus, a step forward in the noise-adapted optimal quantum error mitigation techniques. The effectiveness of the optimized sequences should certainly be notable for other qubit systems where white noise is not the leading type of the environmental noise.
IV.5 Rapid Time Dependent Noise Spectroscopy
@@
In order to capture the time-evolution of the noise around IBM qubits, we performed time-dependent noise spectroscopy experiments on a qubit. To achieve this, ten $T_{1}$ and ten $T_{2}$ measurements were performed consecutively while keeping the experimental parameters unchanged. Note that each run consists of 4000 shots to attain sufficient signal-to-noise ratio which required approximately 10 minutes to complete. The results displayed in Fig.5(a) and (b) demonstrate that we are indeed able to record the rapid changes in $T_{1}$ and $T_{2}$ curves. Moreover, the extracted noise spectra presented in Fig.5(c) indicate that the discontinuous jumps in the dephasing and relaxation data are neither time-correlated nor entirely interdependent. These observations confirm that our noise spectroscopy model is a very effective tool to extract quasi-instantaneous noise profiles.
@@ V Outlook
Owing to the universality of the technique discussed here, we anticipate that this study will help to examine and benchmark various quantum systems in the near future. We expect that the noise characterization methodology discussed here will provide useful insights to improve the engineering aspects of quantum systems. The fidelity of the current methodology can be improved by taking into account advanced gates Baum et al. (2021), or DRAG pulses Werninghaus et al. (2021); Hyyppä et al. (2025), which are optimized to reduce leakage error, instead of square pulses. This should be possible owing to the recent advancements in the filter function formalism, which allows for analytical representation of arbitrary quantum control sequence Soare et al. (2014); Cerfontaine et al. (2021); Chalermpusitarak et al. (2021); Le et al. (2022). Moreover, if this methodology is employed in conjunction with appropriate instrumentation such as Field Programmable Gate Arrays (FPGAs) Goldblatt et al. (2024), then instantaneous error mitigation strategies can potentially be developed.
Finally, it is important to emphasize that the decoherence functional, $\chi(t)$, which depends on the knowledge of the environmental noise spectrum, is intricately linked to the error rates of the various noise operators in the Kraus picture. Since the Kraus formalism is commonly employed in the design and implementation of QEC protocols, having precise knowledge of the decoherence functional opens up new avenues for noise-adapted error correction protocols Biswas et al. (2024). Indeed, QEC protocols can then be tailored to deal with a specific spectral range of noise. With the help of a trained model, time-dependent quantum error correction protocols can become a foreseeable reality.
@@ VI Acknowledgements
B. G. and V. J., contributed equally to this work. S.D. thanks the Indian Institute of Technology, Madras, India, for its seed and initiation grants. The authors acknowledged the use of facilities supported by a grant from the Mphasis F1 Foundation given to the Centre for Quantum Information, Communication, and Computing (CQuICC). Finally, the authors appreciated the use of IBM Quantum for this work. The views expressed were those of the authors and did not reflect the official policy or position of IBM or the IBM Quantum team.
Data Availability
The relevant experimental data and the trained model parameters are available from the authors.
Code Availability
The code for training data generation and the model construction is provided in the public repository.
@@ Appendix A Noise Model: Kraus Operator Picture
where the physical process map $\mathcal{E}$ on the system operator Hilbert space $\mathcal{H}_{S}$ must be completely positive (CP). The complete positivity of the map implies the existence of an operator–sum decomposition for the map Nielsen and Chuang (2011), of the following form
@@
We will now describe two examples of such CPTP maps — also referred to as quantum channels Nielsen and Chuang (2011)— which become important in the context of superconducting qubits. The first is the so called amplitude damping channel, which characterizes the effects due to loss of energy from a quantum system. Specifically, it describes energy dissipation in a two-level system. Let $|0\rangle$ denote the ground state and $|1\rangle$ some excited state of a qubit. Then, the amplitude damping channel denoted as $\mathcal{E}^{AD}=\{E_{0}^{AD},E_{1}^{AD}\}$, is described by the following pair of Kraus operators Nielsen and Chuang (2011).
@@
$=\begin{pmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{pmatrix}$,
@@
$E_{0}^{PD}$ $=\sqrt{p}I=\begin{pmatrix}\sqrt{p}&0\\ 0&\sqrt{p}\end{pmatrix}$,
@@
$\rho$ $=\begin{pmatrix}|\alpha|^{2}&\alpha\beta^{*}\\ \alpha^{*}\beta&|\beta|^{2}\end{pmatrix}$ (8)
$\rho_{noisy}$ $=\mathcal{E}^{AD}\circ\mathcal{E}^{PD}(\rho)$ (9)
$=\mathcal{E}^{PD}\circ\mathcal{E}^{AD}(\rho)$ (10)
$\rho_{noisy}$ $=\begin{pmatrix}1+(|\alpha|^{2}-1)e^{-\Gamma_{1}t}&\alpha\beta^{*}e^{-\frac{\Gamma_{1}}{2}t}e^{-\Gamma_{\phi}t}\\ \alpha^{*}\beta e^{-\frac{\Gamma_{1}}{2}t}e^{-\Gamma_{\phi}t}&|\beta|^{2}e^{-\Gamma_{1}t}\end{pmatrix}$ (11)
@@ Appendix B Neural Network
We used TensorFlow Keras Python module to train our convolutional neural network and to perform the noise spectra predictions from the experimentally obtained decoherence curves. The autoencoder-type network, as described in Table 2 alternates between convolutional and pooling operations to progressively extract features from the input data, and the upsampling restores the data’s original dimensions while enhancing feature representation. The final dense layer aims to produce a linear output appropriate for regression or other continuous output tasks. Our network is trained for approximately 75 epochs, achieving an accuracy loss lower than 4%. The training and validation losses, measured as mean absolute error, are monitored over epochs.
Figure 6: Distribution of fitted $p$ and $T_{\phi}$ in the training dataset
Furthermore, to obtain a high degree of prediction accuracy, it is essential to constrain the training data appropriately. We fit the generated decoherence curves using a stretched exponential function:
$\tilde{C}(t)=e^{-(t/T_{\phi})^{p}}$ (12)
Although, the simulated curves cannot be exactly fitted with Eq.12, this method provides a good measure to filter the training data. SI Figure 6 shows the final distribution of the stretching factor $p$, and of inverse decay rate $T_{\phi}$ in the training dataset. All the experimental curves lie well within the chosen bounds.
Appendix C Dynamical Decoupling Optimization
To determine the optimal number of pulses for practical implementation, we performed a systematic analysis comparing CPMG sequences with a varied number of pulses $(n=4,8,16,32)$. Fig.7 demonstrates the $\%$ coherence improvement as a function of evolution time for these different sequences. Supposing that the gain is relevant only when it occurs at short evolution times ($T_{2}$), it can be postulated that 8 $\pi$-pulse sequences are close to optimal. The $\%$ improvements observed at longer evolution times, specifically for $n=16,32$ are practically ineffective.
Figure 7: $\%$ Coherence Improvement for a randomly chosen qubit as a function of evolution time for CPMG pulse sequences with nn=4,8,16, and 32 $\pi$-pulses each having a width of 48 ns.
Appendix D Classical Optimization vs Machine Learning (ML) based Approach
In order to validate the superiority of our approach we compared the results shown in Figure 2 of the manuscript with the state-of-the-art classical optimizers as depicted in the SI Figure 3. For this purpose, we implemented brute force optimization using the functional form of $S(\omega)$ described in Section 3 of the manuscript. We used the Constrained Optimization BY Linear Approximation (COBYLA) algorithm provided by SciPy. We analyzed 200 test datasets having the same percentage of experimental noise and repeated the procedure for various noise percentages. For the ML-based approach, the error distributions (presented in Figure 2 of the manuscript) have a well-defined maxima; therefore, we extracted the mode value associated with each histogram and considered the standard deviation around the value to generate the error bars. Whereas, the error distribution in the case of classical optimization does not show a clear trend, and thus, we use the mean value and standard deviation around the mean to analyze the data. The comparative plot evidently shows that the ML-based technique outperforms the classical optimization approach even in the case of the data with a low signal-to-noise ratio.
Figure 8: $\%$ error in the predicted $S(\omega)$ for the two distinct methodologies. The histograms at the bottom, show the distribution of the $\%$ errors in the individual predictions for the classical optimization protocol. These histograms are shown for the datasets comprising the lowest and highest experimental noise. The green dashed line represents the mean, while the blue dashed line represents the $\%$ error 1 standard deviation away from the mean value.